OpenStudy (anonymous):

A buoy is made from a steel tube 4meters long and 0.5meters in radius, with a welded one half-sphere at each end. Knowing the weight of the buoy is 3630N *) Find the height of immersion of the buoy if it is immersed in fresh water. The buoyancy force is directed vertically upwards when a body is immersed in a liquid (allowing it to float). This force corresponds to the weight of the liquid displaced. For example, an object immersed in water (with a density of 1000 kg/m3) undergoes a buoyancy of 9810N for each cubic meter of its volume submerged (that is to say 9 • 81m/s2 1000 kg/m3 • 1m3 = 9810N).

OpenStudy (anonymous):

I've calculated the volume of the buoy (volume of cylinder + volume of sphere), which equals 3.66519 cubic meters \[V =(π×〖0.5m〗^2×4m)+ ((4π〖0.5〗^3)/3) \]

OpenStudy (anonymous):

\[Buoy's Buoyancy force = 9810N * 3.66519 = 35955.5139 Newtons\]

OpenStudy (anonymous):

Not sure if useful, but I've calculated the Buoy's mass in kg \[Buoy's Mass = 3630 Newtons (N = mg, N/g=m, 3630N/9.81m/s ^{2}=370.03kg\]

OpenStudy (anonymous):

and from here I'm stuck, I know I should be using integrals but I'm not sure how to go about it to find the immerged height

OpenStudy (anonymous):

will pay for help. lol